If i could please get some help with this i have tried every source but this one please the problem is y-3=2/5(x+1) please help me.
Answer by lwsshak3(11628)

If i could please get some help with this i have tried every source but this one please the problem is y-3=2/5(x+1).
y-3=2/5(x+1)
y-3=(2/5)x+2/5
y=(2/5)x+2/5+3
y=(2/5)x+2/5+15/5
y=(2/5)x+17/5
This is an equation of a straight line with slope=(2/5) and y-intercept=17/5